### v1

@@ 6 Related Work
The evolution of large language models (LLMs) Touvron et al. (2023); Bai et al. (2023); Achiam et al. (2023) is a remarkable milestone while addressing concerns of reliability and safety. The development of the generative pre-trained Transformer (GPT) series by OpenAI progressively expands the frontiers of the field of natural language processing (NLP), such as named entity recognition Mo et al. (2024a); Yang et al. (2022); Mo et al. (2024b) and machine translation Yang et al. (2020b, 2021, a, 2023). To further enhance the adaptability of LLMs to various tasks, instruction-tuning has become a commonly adopted method by utilizing instructions derived from human-curated prompts, feedback, and public benchmarks, which supports the fast development of the domain-specific LLMs, such as complex reasoning Chai et al. (2024); Qin et al. (2024), coding Roziere et al. (2023); Lozhkov et al. (2024), IT operation Guo et al. (2023, 2024), and science Xie et al. (2023). Instruction tuning Wang et al. (2023) is introduced to generate novel tasks and associated instructions from the ground up, offering advantages in terms of performance and cost efficiency. The breadth and variety of the instruction data correlate strongly with how well the models can generalize and handle tasks they have not encountered before.
@@ 7 Conclusion
In this work, we present SEvenLLM, which marks a significant advancement in the use of LLMs for benchmarking, eliciting, and improving cybersecurity incident analy- sis and response abilities in cybersecurity LLMs. By leveraging a meticulously curated bilingual instruction corpus from an extensive collection of cybersecurity texts, SEvenLLM bridges the gap in the availability of effective data for cybersecurity applications. The instruction dataset SEvenLLM-Instruct encompassing 27 well-conceptualized tasks is used to fine-tune SEvenLLM based on the different foundation models (Llama and QWen). The extensive main and analytic experiments performed on a specialized curated cybersecurity benchmark, SEvenLLM-Bench, further corroborate the efficacy of SEvenLLM in improving the analytical capabilities and providing a robust response mechanism against cyber threats.

### v2

@@ 6 Related Work
The evolution of large language models (LLMs) Touvron et al. (2023); Bai et al. (2023); Achiam et al. (2023) is a remarkable milestone while addressing concerns of reliability and safety. The development of the generative pre-trained Transformer (GPT) series by OpenAI progressively expands the frontiers of the field of natural language processing (NLP), such as named entity recognition Mo et al. (2024a); Yang et al. (2022); Mo et al. (2024b), machine translation Yang et al. (2020b, 2021, a, 2023) and text style transfer Wang et al. (2019, 2020); Wu et al. (2020). It even has an impact Lin et al. (2023); Zhai et al. (2024) on some traditional recommendation and advertising methods Wang et al. (2023b); Huang et al. (2022). To further enhance the adaptability of LLMs to various tasks, instruction-tuning has become a commonly adopted method by utilizing instructions derived from human-curated prompts, feedback, and public benchmarks, which supports the fast development of the domain-specific LLMs, such as complex reasoning Chai et al. (2024); Qin et al. (2024), coding Roziere et al. (2023); Lozhkov et al. (2024), IT operation Guo et al. (2023, 2024), and science Xie et al. (2023). Instruction tuning Wang et al. (2023a) is introduced to generate novel tasks and associated instructions from the ground up, offering advantages in terms of performance and cost efficiency. The breadth and variety of the instruction data correlate strongly with how well the models can generalize and handle tasks they have not encountered before.
@@ 7 Conclusion
In this work, we present SEvenLLM, which marks a significant advancement in the use of LLMs for benchmarking, eliciting, and improving cybersecurity incident analy- sis and response abilities in cybersecurity LLMs. By leveraging a meticulously curated bilingual instruction corpus from an extensive collection of cybersecurity texts, SEvenLLM bridges the gap in the availability of effective data for cybersecurity applications. The instruction dataset SEvenLLM-Instruct encompassing 28 well-conceptualized tasks is used to fine-tune SEvenLLM based on the different foundation models (Llama and Qwen). The extensive main and analytic experiments performed on a specialized curated cybersecurity benchmark, SEvenLLM-Bench, further corroborate the efficacy of SEvenLLM in improving the analytical capabilities and providing a robust response mechanism against cyber threats.
Limitations
The primary data source collected focuses on English as the mainstream language for cybersecurity intelligence. Although it has been extended to include Chinese data, the analysis of cybersecurity incidents would benefit from a multilingual capability, which requires further collection and organization. In the future, we will expand SEvenLLM-Instruct to more languages.
Ethical Considerations
In this study, we analyze cyber threat intelligence exclusively derived from publicly available sources provided by security companies. Although some sensitive cybersecurity terminology was used, we strictly avoided involving any personal privacy, trade secrets, or activities that could constitute a threat. Moreover, we ensured that all data handling and usage complied with ethical standards and legal regulations to maintain transparency and integrity in our research.